• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Peter L

    Log entry time 03:17:04 on July23,2005

    Entry number 149107

    keyword=Left HRS IOC reboot

    At 01:55 we needed to reboot the Left HRS IOC. Upon returning we needed to ramp Q1,Q2,Q3, and D1 back to their nominal. We also needed to place the D1 regulator back to ON. The magnets ramped while we were in the hall.